OpenStudy (anonymous):

What is am equation for the line that passes through points (-1,-4) and (1,4)? Write the equation in slope intercept form. a. y=4x b. y=-4x c. y= -1x -4 d. y= -4x + 1

OpenStudy (anonymous):

Can you help?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ -4}})\quad &({\color{red}{ 1}}\quad ,&{\color{blue}{ 4}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (anonymous):

So what? And can you help me with more?

OpenStudy (jdoe0001):

well, find the slope then plug the values in the point-slope form :) then solve for "y"

OpenStudy (anonymous):

I will medal, also are you in connexus?

OpenStudy (anonymous):

I got the answer b.

OpenStudy (jdoe0001):

what did you get for the slope?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Never mind I'll gues can we do another?

OpenStudy (anonymous):

Can you tell me if Im right?

OpenStudy (jdoe0001):

well... you can always post anew \(\bf \bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ -4}})\quad &({\color{red}{ 1}}\quad ,&{\color{blue}{ 4}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ 4}}-{\color{blue}{ -4}}}{{\color{red}{ 1}}-{\color{red}{ -1}}}\implies \cfrac{8}{2}\implies 4\)

OpenStudy (jdoe0001):

so now you know the slope, then just plug it in along with the point coordinates :)

OpenStudy (anonymous):

Ok I got c now am I right?

OpenStudy (jdoe0001):

well... try again.... notice the slope is 4 and the point you could use could be either point so using the 1st point given you'd end up with \(\bf y-{\color{blue}{ -4}}={\color{green}{ 4}}(x-{\color{red}{ -1}}) \)

OpenStudy (anonymous):

Ok so 4y=4x-4 right?

OpenStudy (jdoe0001):

well... \(\bf y-({\color{blue}{ -4}})={\color{green}{ 4}}(x-({\color{red}{ -1}}))\)

OpenStudy (anonymous):

Now I think 4y=4x+1

OpenStudy (anonymous):

Am I close now?

OpenStudy (jdoe0001):

well... - * -4 = +4

OpenStudy (jdoe0001):

\(\bf y-({\color{blue}{ -4}})={\color{green}{ 4}}(x-({\color{red}{ -1}}))\implies y+4=4(x+1) \\ \quad \\ \implies y+4=4x+4\implies y=4x+4-4\implies y=4x\)

OpenStudy (anonymous):

Your very good at this, can I ask one more. It was very clear and explained

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Which of the following equations would graph a line parallel to 3x-y=7? a. y=7x-3 b. y=-3x+7 c. y=-7x + 3 d. y=3x

OpenStudy (jdoe0001):

well if you solve -> 3x-y=7 <- for "y" what would it look like?

OpenStudy (anonymous):

not a clue but I think 4x=-y

OpenStudy (jdoe0001):

try adding "y" from both sides, then subtracting "7" to both sides what would that give you?

OpenStudy (anonymous):

So far 3x=7y is this correct?

OpenStudy (jdoe0001):

well... \(\bf 3x-y=7\qquad {\color{red}{ +y}} \\ \quad \\ 3x\cancel{+y-y}=7+y\implies 3x=7+y\qquad {\color{red}{ -7}} \\ \quad \\ 3x-7=\cancel{7}+y\cancel{-7}\)

OpenStudy (anonymous):

Im sorry but I have did every thing and still have not got any of those answers

OpenStudy (jdoe0001):

well. after solving for "y" what did we end up with?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

And one time 4

OpenStudy (jdoe0001):

wel... not quite.. notice -> \(\bf 3x-y=7\qquad {\color{red}{ +y}} \\ \quad \\ 3x\cancel{+y-y}=7+y\implies 3x=7+y\qquad {\color{red}{ -7}} \\ \quad \\ 3x-7=\cancel{7}+y\cancel{-7}\) what's "y" equals to?